SDSS Trip Report
SDSS Interlock System Work
November 16-19, 1998

John Anderson, Glenn Federwitz, and Angela Prosapio
Fermi National Accelerator Laboratory
Batavia, IL   USA    60510

The following work was completed this week on the 2.5-meter telescope interlock system.  This report was prepared by John Anderson.

PLC Power Supply
The power supply for the interlock PLC was replaced with a larger supply. The new power supply is run off the 110 vac UPS in the interlock rack.

PLC Input and Output Modules
The I/O modules for the interlock PLC were replaced with 32 bit I/O modules. This was done to increase the number of available inputs for the interlock system.

Lightning Protection
A lightning surge suppresser was added to the altitude junction box on the remote I/O link to the rotator junction box.

Building-in-place Switch
A new switch was mounted and wired to detect the building in place over the telescope.

Altitude Motion Interlock Change
A new interlock was installed to allow for altitude motion inside the building. With the building in place switch, the windscreen and telescope are allowed to move up to 18.5 degrees. This should help in stowing the telescope inside the building. Currently this is implemented as a soft interlock only. Steve Bastian will modify the Altitude interlock switch enclosure to add an additional switch. The new switch will be added to the low level interlocks in January.

Building Clear Switch
A rail was added to detect the entire safe range of the building off the telescope.

Telescope Azimuth Alignment Clamp
We were able to work out with Charlie Briegel the interface problems between the interlock system and the MCP. We were able to activate the clamp under MCP control and read the status correctly.

Telescope Fiducial System Cables
Fiducial and bar code reader cables were pulled for the azimuth and altitude axis fiducial systems. The cables will need to be terminated at a later trip.

A new cable track was mounted inside the cone area of the telescope to support the fiducial and bar code cables. More track will need to be ordered and installed at a later date.

Splitter Chassis Bypasses
With the above interlock modifications, I believe there is no need for any splitter chassis interlocks to be bypassed. The only exception may be the hatch doors.

Altitude and Azimuth Interlock Tests
All altitude and azimuth, soft and hard interlock were tested and found to function correctly. The rotator interlocks will be tested in January after the new rotator is installed. A report on the interlocks tested and their activation points will follow next week.

Spectrograph Slithead Control Boxes
Glenn, Alan Uomoto, and Steve Bastian worked out the mounting scheme for the spectrograph slithead door control boxes. Steve will begin mounting the boxes when he has time.

Azimuth Amplifier #1 Failure
The azimuth amplifier #1 voltage comparator board failed this week. This is the same problem identified on October 29, 1998 with IC1 failing. We replaced the board with a board from the spare Alt/Az amplifier. All seems to be working correctly.

The board in the spare Alt/Az amplifier has been repaired and modified with transient supressors across the Glentek +/- 15 vdc power supply lines and a series resistor to IC1. If the azimuth amplifier fails again, replace the board with the modified board.

Questions regarding this trip report should be addressed to John Anderson at (630) 840-4973.

Last updated 11/19/98.